This ellipse is centered at the origin and has a horizontal axis of length 16 and a vertical axis of length 24. What is its equation?
Answer by lwsshak3(11628)

This ellipse is centered at the origin and has a horizontal axis of length 16 and a vertical axis of length 24. What is its equation.
**
Given equation is that of an ellipse with vertical major axis of the standard form:
(x-h)^2/b^2+(y-k)^2/a^2=1,a>b, with (h,k) being the (x,y) coordinates of the center.
..
center: (0,0)
..
length of vertical major axis=24=2a
a=12
a^2=144
..
length of minor axis=16=2b
b=8
b^2=64
..
Equation of ellipse:
(x-0)^2/64+(y-0)^2/144=1
=x^2/64+y^2/144=1